Looking for some help setting up and solving this problem-Thanks!: One number is 4 times another. If the sum of their reciprocals is 5/36 , find the two numbers.
Answer by stanbon(48558)

Looking for some help setting up and solving this problem-Thanks!: One number is 4 times another. If the sum of their reciprocals is 5/36 , find the two numbers.
------------
Let one of the numbers be "x"; Let the other be "4x".
EQUATION:
1/x + 1/4x = 5/36
Multiply thru by 36x to get:
36 + 9 = 5x
45 = 5x
x=9
============
Cheers,
Stan H.